# TMT: Tri-Modal Translation between Speech, Image, and Text by Processing Different Modalities as Different Languages

### v1

@@ Abstract
The capability to jointly process multi-modal information is becoming an essential task. However, the limited number of paired multi-modal data and the large computational requirements in multi-modal learning hinder the development. We propose a novel Tri-Modal Translation (TMT) model that translates between arbitrary modalities spanning speech, image, and text. We introduce a novel viewpoint, where we interpret different modalities as different languages, and treat multi-modal translation as a well-established machine translation problem. To this end, we tokenize speech and image data into discrete tokens, which provide a unified interface across modalities and significantly decrease the computational cost. In the proposed TMT, a multi-modal encoder-decoder conducts the core translation, whereas modality-specific processing is conducted only within the tokenization and detokenization stages. We evaluate the proposed TMT on all six modality translation tasks. TMT outperforms single model counterparts consistently, demonstrating that unifying tasks is beneficial not only for practicality but also for performance.
1 Introduction
$*{}^{*}$qual contribution.
@@ 1 Introduction
In search of artificial general intelligence, seamlessly processing and representing multi-modal information is a crucial prerequisite. Leveraging numerous efforts on the investigation of diverse multi-modal work, systems capable of processing audio-visual Chen et al. (2021); Hong et al. (2023), audio-text Yang et al. (2023); Wu et al. (2024), and text-visual Ramesh et al. (2021); Li et al. (2023b) modalities are actively emerging. Furthermore, developing a unified model that processes tri-modalities (i.e., audio, visual, and text) is drawing attention given its potential for advancing general intelligence Wu et al. (2023); Han et al. (2023).
@@
However, building such a unified model remains challenging due to multifaceted factors. The distinct nature of different modalities intrinsically burdens the multi-modal systems to include two roles: processing modality-specifics and translating between different modalities. Consequently, the computational load required to train multi-modal models usually exceeds that of uni-modal models, attributed to the effort necessary to do two things simultaneously [^1] We believe this computational burden is one of the most impactful bottlenecks that hinder large multi-modal models, compared to Large Language Models (LLMs), where one can scale the model size and efficiently train it with text-only data.. Furthermore, the necessity of paired multi-modal data is another obstacle because the amount of such data (e.g., speech-image-text) is less abundant than uni-modal data.
@@
In this paper, we propose a novel Tri-Modal Translation (TMT) framework that freely translates between different modalities (i.e., speech, image, and text) using a single model. Specifically, we present a novel viewpoint on multi-modal modeling by interpreting different modalities as different languages. We split the Multi-Modal Translation (MMT) problem into modality-specific processings and an integrated translation, where the latter is treated as the Neural Machine Translation (NMT). To this end, we discretize all modalities into tokens in advance [^2] Image and speech are represented as sequences of integers. by employing pre-trained modality-specific tokenizers. This enables us to leverage the expertise from each field, discarding modality-specifics from the core translation modeling.
@@
The proposed TMT framework and viewpoint introduce several benefits. First, the proposed system is “purely multi-modal” by uniformly interfacing all modalities together different to the method of mapping different modalities into the text space of LLM Wu et al. (2023), which we believe is the future direction of building multi-modal LLMs. Second, through discretization, the computational burden in modeling speech and image is significantly relaxed, specifically by more than $99$Chang et al. (2023b); Kim et al. (2024a) in the aspect of required bits. Finally, we can employ unlimited training data through Back Translation (BT) Sennrich et al. (2016), a well-known data augmentation strategy in NMT333 BT typically translates target monolingual data into a source language to construct source-to-target parallel data.. Therefore, even utilizing uni-modal data Deng et al. (2009); Kahn et al. (2020); Fan et al. (2019) in building TMT becomes possible. With these advantages, we can efficiently scale up multi-modal training, combining computational efficiency with large-scale data.
@@
The contributions of this paper can be summarized as follows: 1) To the best of our knowledge, this work is the first to explore translations between tri-modalities, image, speech, and text, by discretizing all modalities. 2) We provide a novel interpretation of modalities as languages. By regarding MMT as NMT, we show that target uni-modal corpora can be also employed through back translation. 3) The proposed TMT encompasses direct speech-to-image synthesis and image-to-speech captioning, both of which have not well been addressed in the previous literature. 4) We show that the proposed TMT can be efficiently trained like text-only systems via extensive experiments on six MMT tasks. Project page is available on github.com/ms-dot-k/TMT.
@@
2 Related Work
NMT aims to transform a source language sentence into a target language Stahlberg (2020). With advanced network architectures Kenton and Toutanova (2019); Lewis et al. (2020) and training methodologies, including denoising pre-training Ho et al. (2020) and BT Sennrich et al. (2016), the field has rapidly developed. In this work, we interpret discretized speech and images as different languages so that TMT can be addressed using the format of NMT. In this viewpoint, NMT techniques, back translation and network architecture, are incorporated into TMT.
MMT. Developing models capable of performing multiple MMT tasks is a popular research topic with diverse groups of researchers’ investigation Kim et al. (2023); Ao et al. (2022); Girdhar et al. (2023). MUGEN Hayes et al. (2022) explored the discretization of video, audio, and text obtained from synthetic game data, and successfully trained multiple MMT models capable of a single MMT task. VoxtLM Maiti et al. (2024) proposed a decoder-only architecture to train a model that performs four MMT tasks ($2\times 2$) spanning speech and text modalities. They discretized speech using a pre-trained self-supervised feature and treated discretized speech tokens in the same manner as text tokens within a combined vocabulary set. SEED Ge et al. (2023a, b) proposed an image tokenizer using contrastive learning between image and text. With the learned image tokens, they finetuned an LLM spanning text and image modalities, covering two MMT tasks, image captioning and text-to-image synthesis. Different to VoxtLM and SEED, TMT is not restricted to bi-modal translation and can translate between tri-modalities.
There also exist a few concurrent works aiming to handle tri-modalities. NExT-GPT Wu et al. (2023) tries to adapt different modalities into the text space of a pre-trained LLM by using several modality-specific encoders and adaptors, where the overall performance depends on the performance of the LLM. Distinct from NExT-GPT, our TMT is “born multi-modal” by learning from multi-modal data instead of projecting different modalities’ representation into that of a text LLM. Gemini Team et al. (2023) is a recently proposed multi-modal LLM. Different to Gemini, the proposed TMT can directly synthesize speech. Most significantly, different to the existing work, TMT treats different modalities as different languages by interpreting discretized image and speech data equivalent to text tokens. Hence, modality specifics are separated and all six MMT tasks are jointly modeled with a unified interface with fully shared parameters.
3 Method
[…]
3.1 Individual tokenization of each modality
@@
Discretized representations have demonstrated promising results in various research fields. In image processing, Vector Quantization (VQ) Van Den Oord et al. (2017); Esser et al. (2021); Ge et al. (2023a, b) has successfully expressed images using discrete tokens. In speech processing, VQ Défossez et al. (2022); Wang et al. (2023) and Self-Supervised Learning (SSL) Baevski et al. (2020); Hsu et al. (2021a); Chen et al. (2022) model-based discrete tokens Lakhotia et al. (2021) are showing remarkable performances across different tasks (e.g., speech translation Lee et al. (2022b); Popuri et al. (2022); Choi et al. (2023) and synthesis Polyak et al. (2021); Hayashi and Watanabe (2020)).
@@
Motivated by the recent success, we tokenize image, speech, and text modalities. Specifically, as shown in Fig. 1a, the raw image is tokenized into image tokens $u_{i}\in\mathbb{N}^{L_{i}}$ by using SEED-2 Ge et al. (2023b) tokenizer, where each image is compressed into 32 tokens (i.e., $L_{i}=32$). Speech is also tokenized into discrete values $u_{s}\in\mathbb{N}^{L_{s}}$ by clustering the features of a HuBERT Hsu et al. (2021a), an SSL speech model, whose output is 50Hz. Following the practices in Lakhotia et al. (2021), we merge duplicated tokens so that the granularity of speech tokens (i.e., $L_{s}$) is less than 50Hz. For the text input, we employ the BERT tokenizer Schuster and Nakajima (2012); Kenton and Toutanova (2019) to derive text tokens $u_{t}\in\mathbb{N}^{L_{t}}$.
Advantages of tokenizing tri-modalities. Tokenizing all three modalities brings two main advantages. Firstly, the MMT problem now has a unified interface. Modality specifics are processed mostly within tokenization and detokenization process.
@@
Table 1: Data size (bits) comparisons between raw data and tokens. The values are based on the 16kHz 16bits for raw speech waveforms, 224$\times$224 for raw RGB image, 80 filterbanks with 10ms shift for mel-spectrogram. $L$ is the length (sec) of the audio. For text, one character is assigned 8 bits. $S$ refers to the length of the characters.
The other key advantage is efficiency Chang et al. (2023b); Park et al. (2023). Processing raw image and audio is burdensome due to the necessity of large-scale data storage and computational resources Kim et al. (2024b). Discretizing speech and image reduces the bits required to less than 0.2% and 0.04% as shown in Table 1 and appendix F. Therefore, by tokenizing all modalities before training, we can significantly increase the computational efficiency for both storage and computing memory.
3.2 Integrated modeling of tri-modalities
As shown in Fig. 1b, we leverage a multi-modal encoder-decoder architecture Vaswani et al. (2017); Lewis et al. (2020); Liu et al. (2020), widely used in the NMT literature as the backbone architecture. Therefore, all parameters are shared between all six MMT tasks. A popular auto-regressive training scheme is employed. To inform both the encoder and the decoder about input and output modalities, we add a modal-type embedding to the embedded features after the token embedding layer, similar to language embeddings Conneau and Lample (2019). Then, the losses for each MMT are summed and backpropagated. Formally put, the objective function of TMT training can be expressed as:
@@ 3.2 Integrated modeling of tri-modalities
3.3 MMT tasks covered by TMT
TMT incorporates all six MMT tasks (${}_{3}\mathbf{P}_{2}$) spanning different combinations of input and output modalities. Related works for each MMT task can be found in appendix E.
Image Captioning Xu et al. (2015) is a task of describing input image content using natural language. A comprehensive understanding of the input image and the naturalness of the output text are important.
@@ 3.3 MMT tasks covered by TMT
Image normal-$\rightarrow$ Text Single MMT 28.9 27.3 53.6 101.9 20.5 21.8 24.3 49.7 67.7 19.1
@@
Image normal-$\rightarrow$ Speech Single MMT 22.5 21.5 47.1 69.1 15.0 16.9 20.0 45.6 44.7 13.5
@@
Table 2: Comparisons between single MMT systems and TMT on the image and speech-driven captioning tasks. All models are trained on combined COCO and Flickr8k train sets.
@@
Text normal-$\rightarrow$ Image Single MMT 68.8 - 69.3 -
@@
Speech normal-$\rightarrow$ Image Single MMT 57.5 - 58.5 -
@@
Speech normal-$\rightarrow$ Text Single MMT - 6.9 - 6.8
@@
Text normal-$\rightarrow$ Speech Single MMT - 16.0 - 9.7
@@
Table 3: Comparisons between single MMT systems and TMT on four MMT tasks. All models are trained on combined COCO and Flickr8k train sets.
Image-to-Speech Captioning Hsu et al. (2021b) is a task that directly synthesizes a spoken description of the input image. This task is used for the image-to-speech translation evaluation of TMT. It is regarded as a challenging task due to insufficient supervision and the complex nature of speech.
Text-driven Image Synthesis Ramesh et al. (2021, 2022); Saharia et al. (2022); Chang et al. (2023a) aims to synthesize an image that matches a given text description. Both the naturalness of the generated image and the correctness of description are important.
Speech-driven Image Synthesis generates an image from a spoken description Li et al. (2020); Wang et al. (2020). The task is equal to solving an automatic speech recognition (ASR) and text-driven image synthesis at once without the intermediate text representation. Similar to the relationship between image captioning and image-to-speech captioning, further challenges exist, mainly due to the absence of text modality. This task remains crucial for those languages that do not have writing systems Lee et al. (2022b), where text-based systems cannot be applied. Note that this paper is the first to synthesize high-resolution images of 768 $\times$ 768 from speech.
ASR is a well-established task that generates text transcriptions from an input speech Amodei et al. (2016); Prabhavalkar et al. (2023). We use this task for speech-to-text translation of TMT.
Text-to-Speech Synthesis (TTS) generates spoken content corresponding to the given text transcription Wang et al. (2017); Tan (2023).
3.4 Back Translation (BT)
[…]
4 Experiments
4.1 Datasets
The training data comprises Conceptual Captions 3M (CC3M), Conceptual Captions 12M (CC12M) Sharma et al. (2018); Changpinyo et al. (2021), COCO Lin et al. (2014), SpokenCOCO Hsu et al. (2021b), Flickr8k Hodosh et al. (2013), and Flickr8kAudio Harwath and Glass (2015). For COCO and Flickr8k, we employ the original corpora for image-text pairs and then further employ SpokenCOCO and Flickr8kAudio, their recorded speech version, to comprise audio-text-image tri-modal pairs. For CC3M and CC12M, VITS Kim et al. (2021a), a TTS model trained on VCTK Yamagishi et al. (2019), is employed to synthesize speech from random speakers to compile audio-text-image pairs. The evaluation is performed on the test split of COCO and Flickr8k after finetuning. The popular Karpathy splits Karpathy and Fei-Fei (2015) are employed for COCO and Flickr8k.
4.2 Metrics

### v2

@@ Abstract
The capability to jointly process multi-modal information is becoming essential. However, the development of multi-modal learning is hindered by the substantial computational requirements and the limited availability of paired multi-modal data. We propose a novel Tri-Modal Translation (TMT) model that translates between arbitrary modalities spanning speech, image, and text. We introduce a simple yet efficient and effective approach, treating speech and image modalities as discrete text modality and approaching multi-modal translation as a well-established machine translation problem. To this end, we tokenize speech and image data into discrete tokens, resulting in a significant reduction in computational cost. Furthermore, by incorporating back translation into multi-modal translation, unpaired data can also be utilized for training. TMT can perform six modality translation tasks and consistently outperforms its single-model counterparts. TMT significantly reduces the required data size (in bits) for training, to approximately 0.2% for speech data and 0.04% for image data, respectively.
Index Terms:
Tri-modal translation, Text-to-speech, Text-to-image, Image-to-text, Image-to-speech, Speech-to-text, Speech-to-image.
I Introduction
@@
In search of artificial general intelligence, seamlessly processing and representing multi-modal information is a crucial prerequisite. Leveraging numerous efforts on the investigation of diverse multi-modal work [1, 2], systems capable of processing audio-visual [3, 4, 5], audio-text [6, 7], and text-visual [8, 9] modalities are actively emerging. Furthermore, developing a unified model that processes tri-modalities (i.e., audio, visual, and text) is drawing attention given its potential for advancing general intelligence [10, 11].
@@
However, building such a unified model remains challenging due to multifaceted factors. Essentially, the data size (in bits) of speech and image modalities exceeds that of text by a significant margin. Furthermore, the distinct nature of different modalities intrinsically burdens the multi-modal systems to include two roles: processing modality-specifics and translating between different modalities. Consequently, the computational load required to train multi-modal models usually exceeds that of uni-modal models, attributed to the large data size and the effort necessary to do two roles simultaneously [^1] We believe this computational burden is one of the most impactful bottlenecks that hinder large multi-modal models, compared to Large Language Models (LLMs), where one can scale the model size and efficiently train it with text-only data.. Furthermore, the necessity of paired multi-modal data is another obstacle because the amount of such data (e.g., speech-image-text) is less abundant than uni-modal data.
@@
In this paper, we propose a novel Tri-Modal Translation (TMT) framework that freely translates between different modalities (i.e., speech, image, and text) using a single model. Specifically, we present a novel viewpoint on multi-modal modeling by interpreting different modalities as text modality with different languages. We split the Multi-Modal Translation (MMT) problem into modality-specific processings and an integrated translation, where the latter is treated as the Neural Machine Translation (NMT). To this end, we discretize all modalities into tokens in advance [^2] Image and speech are represented as sequences of integers. by employing pre-trained modality-specific tokenizers. This enables us to leverage the expertise from each field, discarding modality-specifics from the core translation modeling.
@@
The proposed TMT framework and viewpoint introduce several benefits. First, the proposed system is “purely multi-modal” by uniformly interfacing all modalities together different to the method of mapping different modalities into the text space of LLM [10], which we believe is the future direction of building multi-modal LLMs. Second, through discretization, the computational burden in modeling speech and image is significantly relaxed, specifically by more than $99$[12, 13] in the aspect of required bits. Finally, we can employ unpaired training data through Back Translation (BT) [14], a well-known data augmentation strategy in NMT333 BT typically translates target monolingual data into a source language to construct source-to-target parallel data.. Therefore, even utilizing extra uni-modal data [15, 16, 17] in building TMT becomes possible. With the advantages of computational efficiency and the accessibility of unpaired data, multi-modal training holds the potential for scalable expansion.
@@
The contributions of this paper can be summarized as follows: 1) To the best of our knowledge, this work is the first to explore translations between discretized (i.e., tokenized) tri-modalities of image, speech, and text. By handling multi-modal data with discretized tokens, we can greatly improve data efficiency, where the required bits for images are reduced to 0.035% of their original size and 0.2% for speech. 2) We treat the discretized modalities as text modality and regard them akin to distinct languages. By applying back translation, an NMT technique, we show that target uni-modal corpora can be also employed for MMT. 3) The proposed TMT encompasses direct speech-to-image synthesis and image-to-speech captioning, both of which have not been well addressed in the previous literature. 4) We show that the proposed TMT can be efficiently trained like text-only systems via extensive experiments and analysis on six MMT tasks. The pre-trained model, code, and qualitative results are available on github.com/ms-dot-k/TMT.
II Related Work
II-A Multi-Modal Translation (MMT)
Developing models capable of performing multiple MMT tasks is a popular research topic with diverse groups of researchers’ investigation [18, 19, 20]. MUGEN [21] explored the discretization of video, audio, and text obtained from synthetic game data, and successfully trained multiple MMT models capable of a single MMT task. VoxtLM [22] proposed a decoder-only architecture to train a model that performs four MMT tasks ($2\times 2$) spanning speech and text modalities. They discretized speech using a pre-trained self-supervised feature and treated discretized speech tokens in the same manner as text tokens within a combined vocabulary set. SEED [23, 24] proposed an image tokenizer using contrastive learning between image and text. With the learned image tokens, they finetuned an LLM spanning text and image modalities, covering two MMT tasks, image captioning and text-to-image synthesis. Different to VoxtLM and SEED, TMT is not restricted to bi-modal translation and can translate between tri-modalities.
There also exist a few concurrent works aiming to handle tri-modalities. NExT-GPT [10] tries to adapt different modalities into the text space of a pre-trained LLM by using several modality-specific encoders and adaptors, where the overall performance depends on the performance of the LLM. Distinct from NExT-GPT, our TMT is “born multi-modal” by learning from multi-modal data instead of projecting different modalities’ representation into that of a text LLM. Gemini [25] is a recently proposed multi-modal LLM. Different from Gemini, the proposed TMT can directly synthesize speech. Most significantly, different from the existing work, TMT interprets both discretized image and speech data as text tokens. Hence, the computational efficiency in handling multi-modal data is greatly improved. TMT successfully models all six MMT tasks with a unified interface with fully shared parameters.
II-B Image Captioning
Image Captioning [26] is a task to describe input image content in natural language. It has shown remarkable progress with the development of computer vision and natural language processing technologies. Recent methods showed that vision-language pre-training can significantly improve image captioning performances over training the model from scratch [27]. For example, [28, 29] proposed cross-modal contrastive learning, and [30, 31, 32, 33] proposed multi-task pre-training which contains such as image-text matching, masked language modeling, and contrastive learning. Basically, the task can be regarded as an MMT problem of image-to-text translation.
Similar yet distinct from image captioning, image-to-speech captioning [34] aims to directly generate a spoken description of the input image. As the task does not go through with text modality, it is regarded as a challenging problem due to insufficient supervision and the complex nature of speech. Recent methods tried to mitigate the challenge by representing speech in pseudo-text format and showed promising results [34, 35, 13]. This task falls in text-to-speech translation.
II-C Language-driven Image Synthesis
With the great development of visual codebook representation [36, 37] and diffusion [38, 39], image generation has become more and more realistic. Especially, [8, 40, 41, 42, 43] showed that generative models can now freely create images based on text descriptions by conditioning the text representations into the decoding process, such as diffusion process. Recently, [44] proposed a LLM-based text-to-image synthesis model that first trains with image captioning to enable the LLM to understand image representations. Different from them, the proposed TMT can perform 6 different tasks with a single trained model. Moreover, the proposed method handles the data efficiently through discretized image and speech tokens.
Different from text-based image synthesis, direct speech-to-image synthesis [45, 46] has not well been addressed. Since many languages have no writing systems [47], text-based systems cannot be applied to such languages. Hence, the demand for technology enabling spoken language-driven image synthesis, the speech-to-image translation, is expected to increase.
II-D Neural Machine Translation (NMT)
NMT aims to transform a source language sentence into a target language [48, 49]. With the advent of advanced network architectures [50, 51] and innovative training methodologies, such as denoising pre-training [38] and BT [14], the field has rapidly evolved. Furthermore, NMT has expanded to encompass multi-modal systems, including speech-to-text machine translation [52, 53], speech-to-speech machine translation [54, 55], and text-image machine translation [56, 57]. Among them, BT aims to produce source-target language pairs by performing translation onto the monolingual target language data using an intermediate developing multilingual model. It follows the following steps in general: 1) train an intermediate model with supervised data. 2) produce the source-target language pairs by applying the intermediate model onto the target language monolingual data (i.e., applying BT). 3) by using both supervised data and back-translated data, further train the intermediate model to produce the final model. In this work, we treat discretized speech and images as pseudo-text so that TMT can be efficiently addressed using the format of NMT. In this viewpoint, NMT techniques, back translation and network architecture, are incorporated in TMT.
@@ II-D Neural Machine Translation (NMT)
III Method
Fig. 1 illustrates the overall scheme of the proposed TMT framework. Let $x_{i}$ be an image, $x_{s}$ be speech describing the image, and $x_{t}$ be the transcription of the speech. The main objective is to build a single model that freely translates between different modalities (i.e., $x_{t}$, $x_{s}$, and $x_{i}$). For this goal, we first tokenize image, speech, and text data into discretized tokens (Fig. 1a and §$\S$§III-A). The vocabulary set of TMT is a union set of three sets of vocabularies. Then we treat discretized tokens from different modalities as different languages and train the TMT model like training an NMT model (§$\S$§III-B). The proposed TMT can translate between tri-modalities, covering all six MMT tasks (§$\S$§III-C).
III-A Individual tokenization of each modality
@@
Discretized representations have demonstrated promising results in various research fields. In image processing, Vector Quantization (VQ) [36, 37, 23, 24] has successfully expressed images using discrete tokens. In speech processing, VQ [58, 59] and Self-Supervised Learning (SSL) [60, 61, 62] model-based discrete tokens [63] are showing remarkable performances across different tasks (e.g., speech translation [47, 64, 65] and synthesis [66, 67]).
@@
Motivated by the recent success, we tokenize image, speech, and text modalities. Specifically, as shown in Fig. 1a, the raw image is tokenized into image tokens $u_{i}\in\mathbb{N}^{L_{i}}$ by using SEED-2 [24] tokenizer, where each image is compressed into 32 tokens (i.e., $L_{i}=32$). Speech is also tokenized into discrete values $u_{s}\in\mathbb{N}^{L_{s}}$ by clustering the features of a HuBERT [61], an SSL speech model, whose output is 50Hz. Following the practices in  [63], we merge duplicated tokens so that the granularity of speech tokens (i.e., $L_{s}$) is less than 50Hz. For the text input, we employ the BERT tokenizer [68, 50] to derive text tokens $u_{t}\in\mathbb{N}^{L_{t}}$.
Advantages of tokenizing tri-modalities. Tokenizing all three modalities brings two main advantages. Firstly, the MMT problem now has a unified interface. Modality specifics are processed mostly within tokenization and detokenization process. The other key advantage is efficiency [12, 69]. Processing raw image and audio is burdensome due to the necessity of large-scale data storage and computational resources [70]. Discretizing speech and image reduces the bits required to 0.2% and 0.035%, which are similar levels to text data. Table I shows the data size comparisons in bits between different data types of each modality. By employing speech tokens rather than directly using raw audio or its Mel-spectrogram, we significantly reduce the data size to less than 0.2% compared to other formats. Consequently, 1-second speech tokens contain a similar number of bits as text comprising approximately 50 characters. Furthermore, by utilizing image tokens instead of raw images, only 0.035% of the bits are needed to represent the original image content. This substantial reduction equates one image to the text length of 52 characters. Therefore, by representing and employing the tokenized data instead of directly utilizing their raw format, we can efficiently train our TMT like training a text-only system with reduced data storage requirements and computing resources.
TABLE I: Data size (bits) comparisons between raw data and tokens. The values are based on the 16kHz 16bits for raw speech waveforms, 224$\times$224 for raw RGB image, 80 filterbanks with 10ms shift for mel-spectrogram. $L$ is the length (sec) of the audio. For text, one character is assigned 8 bits. $S$ refers to the length of the characters.
@@
III-B Integrated modeling of tri-modalities
As shown in Fig. 1b, we leverage a multi-modal encoder-decoder architecture [71, 51, 72], widely used in the NMT literature as the backbone architecture. Therefore, all parameters are shared between all six MMT tasks. A popular auto-regressive training scheme is employed. To inform both the encoder and the decoder about input and output modalities, we add a modal-type embedding to the embedded features after the token embedding layer, similar to language embeddings [73]. Then, the losses for each MMT are summed and backpropagated. Formally put, the objective function of TMT training can be expressed as:
@@ III-B Integrated modeling of tri-modalities
III-C MMT tasks covered by TMT
TMT incorporates all six MMT tasks (${}_{3}\mathbf{P}_{2}$) spanning different combinations of input and output modalities.
Image Captioning [26] is a task of describing input image content using natural language. A comprehensive understanding of the input image and the accuracy and completeness of the output text are important.
Image-to-Speech Captioning [34] is a task that directly synthesizes a spoken description of the input image. This task is used for the image-to-speech translation evaluation of TMT. It is regarded as a challenging task due to inherent complex nature of speech where speech contains not only linguistic information but also diverse factors of acoustic noise, different tones, delivery fast, etc.
Text-driven Image Synthesis [8, 40, 41, 42] aims to synthesize an image that matches a given text description. Both the naturalness of the generated image and its correctness to the input description are important.
Speech-driven Image Synthesis generates an image from a spoken description [45, 46]. The task is equal to solving an automatic speech recognition (ASR) and text-driven image synthesis at once without the intermediate text representation. Similar to the relationship between image captioning and image-to-speech captioning, further challenges exist, mainly due to the absence of text modality. This task remains crucial for those languages that do not have writing systems [47], where text-based systems cannot be applied. Note that this paper is the first to synthesize high-resolution images of 768 $\times$ 768 from speech.
ASR is a well-established task that generates text transcriptions from an input speech [74, 75]. We use this task for speech-to-text translation of TMT.
Text-to-Speech Synthesis (TTS) generates speech corresponding to the given text transcription [76, 77].
III-D Back Translation (BT)
We investigate whether applying BT in the MMT problem based on our interpretation can help improve the performance even if utilizing uni-modal databases. To this end, TMT model is first pre-trained with approximately 10M paired tri-modal data (detailed in §$\S$§IV-A) to produce an intermediate model. Then, with the intermediate model, we apply BT on a uni-modal image dataset, ImageNet [15], to produce pseudo source modalities of speech and text. This results in 2.6M text-image and speech-image data. We also apply BT on a speech-text bi-modal dataset, CommonVoice [78], and produce additional 2M image-speech and image-text data. By merging the pseudo data on top of the original tri-modal data pairs, we further train the intermediate TMT model and derive the final TMT model. Please note that, we employ the intermediate TMT model trained on tri-modal paired data to perform BT, instead of using other image captioning or ASR models.
TABLE II: Comparisons between single MMT systems and TMT on the image and speech-driven captioning tasks.
@@ III-D Back Translation (BT)
Image $\rightarrow$ Text Single MMT 28.9 27.3 53.6 101.9 20.5 21.8 24.3 49.7 67.7 19.1
@@
Image $\rightarrow$ Speech Single MMT 22.5 21.5 47.1 69.1 15.0 16.9 20.0 45.6 44.7 13.5
@@
TABLE III: Comparisons between single MMT systems and TMT on four MMT tasks.
@@
Text $\rightarrow$ Image Single MMT 68.8 - 69.3 -
@@
Speech $\rightarrow$ Image Single MMT 57.5 - 58.5 -
@@
Speech $\rightarrow$ Text Single MMT - 6.9 - 6.8
@@
Text $\rightarrow$ Speech Single MMT - 16.0 - 9.7
@@
IV Experiments
IV-A Datasets
The training data comprises Conceptual Captions 3M (CC3M), Conceptual Captions 12M (CC12M) [79, 80], COCO [81], SpokenCOCO [34], Flickr8k [82], and Flickr8kAudio [83].
CC3M is an image-text paired dataset that consists of about 3.3M paired instances. Different from the curated style of other image caption annotations, the dataset is collected from the web through an automatic pipeline. Therefore, the image-text pairs represent a wider variety of styles.
CC12M is an image-text paired dataset with about 12M instances, which is adequate for pretraining vision-language systems. The dataset is similarly collected as CC3M with relaxed constraints during its automatic pipeline to further increase the data size.
COCO is a popular dataset for various computer vision tasks. The data provides images with diverse annotations including object segmentations, object bounding boxes, object classes, key points, and text descriptions. We employed the images and corresponding text descriptions to train image-to-text and text-to-image tasks of TMT.
SpokenCOCO is the speech version of COCO dataset. The dataset is collected by recording human speakers reading the text annotations of COCO. By merging SpokenCOCO with COCO, we can build speech-image-text tri-modal paired datasets which makes it possible to train all 6 tasks of TMT.
Flickr8k is a famous image captioning dataset contains 8K images with five different captions for each image. The images are collected from six different Flickr groups to contain diverse scenes and situations.
Flickr8kAudio is the speech version of the Flickr8k dataset, similar to the relationship between COCO and SpokenCOCO. The dataset is constructed by recording the speech from 183 speakers. By merging Flickr8kAudio with Flickr8k dataset, we can build speech-image-text tri-modal paired datasets.
For CC3M and CC12M, VITS [84], a TTS model trained on VCTK [85], is employed to synthesize speech from random speakers to compile audio-text-image pairs. The evaluation is performed on the test split of COCO and Flickr8k after finetuning. The popular Karpathy splits [86] are employed for COCO and Flickr8k.
IV-B Metrics